# Statistically Valid Post-Deployment Monitoring Should Be Standard for AI-Based Digital Health

### v1

@@ 6 Alternative Views
Several well–established research threads could serve as alternatives to the post-deployment monitoring problem. We examine the pos and cons of the top three alternatives to our proposed approach:continual learning, Bayesian change–point detection, and conformal-prediction–based monitoring.
@@
or lifelong learning algorithms update model parameters online to accommodate non-stationary data distributions ring1994continual ; parisi2019continual ; chen2018lifelong . Advantages include low latency adaptation as well as theoretical guarantees. Models can react to drift on the very next batch, which is attractive when label feedback is cheap. Additionally, PAC-Bayesian alquier2016regret or regret bounds are available for certain online update rules. However, this comes with several limitations: label requirement, auditability and traceability and silent failure. State-of-the-art continual learners still rely on frequent ground-truth labels to avoid catastrophic forgetting rolnick2019experience . In clinical applications, those labels are costly or delayed. Regulatory guidelines (e.g. FDA SaMD) require reproducible model versions. Online updates create a moving target that complicates version control, performance analysis, and root-cause analysis fda2017samed . Finally, without an external test, online updates can chase noisy fluctuations and cause silent accuracy drops mallick2022matchmaker , which in turn may widen performance gaps for under-represented sub-groups yang2023subpopshift .
@@
offers a probabilistic approach by maintaining a posterior distribution over run lengths and updating this belief as new data stream in adams2007bayesian ; fearnhead2007line . This method provides coherent uncertainty quantification, allowing drift detection systems to trigger alarms based on posterior probabilities. It is also sequentially efficient: when using conjugate-exponential models, updates can be computed in constant amortized time per observation. However, BOCPD relies heavily on calibrated priors, which are rarely available or reliable in clinical contexts. Furthermore, applying BOCPD to high-dimensional data often requires approximate inference via particle filters or sequential Monte Carlo methods, which may be computationally infeasible for hospital-scale EHR feeds. Importantly, while BOCPD indicates when a change occurred, it does not identify where the change took place—limiting its usefulness in root-cause analysis and mitigation planning.
@@
offer a distribution-free framework for post-deployment monitoring. These methods maintain finite-sample validity under the assumption of exchangeability and have been extended to detect drift by monitoring conformal $p$-values or exchangeability martingales vovk2022algorithmic ; ho2016conformal ; johansson2022conformal . They are attractive in that they do not require strong parametric assumptions, and variants such as Mondrian or conditional conformal predictors can operate effectively using weak labels. Still, this approach has notable weaknesses. The assumption of exchangeability is fragile in real-world settings like healthcare, where temporal, site-level, and treatment-based correlations are pervasive and violate i.i.d. conditions barber2021limits . Furthermore, the signal provided by conformal methods is often blunt: they flag distributional shift only after prediction sets inflate, and do not distinguish between covariate and concept drift, nor do they identify the specific subgroups affected. As such, they tend to be reactive rather than proactive—triggering alarms after performance degrades, rather than before.
@@
Two-sample hypothesis testing uniquely combines rigorous error control, label efficiency, and interpretability. It provides explicit $\alpha$-level guarantees on type I error and power-based control of type II error. Unlike conformal methods, label-free tests such as kernel MMD borgwardt2006integrating compare input distributions directly and highlight the feature regions driving the difference, providing regulator-friendly, interpretable evidence of shift. Detection is also modular: once drift is identified, retraining, recalibration, or continual learning can follow as appropriate. In contrast, continual learning, Bayesian monitoring, and conformal prediction each address part of the monitoring problem but fall short on one or more axes—supervision cost, statistical guarantees, or auditability. Hypothesis testing, by covering all three, is the most robust and regulator-ready foundation for monitoring AI systems in healthcare.
7 Conclusion and Future Directions
@@
Appendix A Performance Metrics
@@
Appendix B Choosing A Statistical Hypothesis Test and Heuristics
@@
In addition to formal test statistics, our framework includes heuristics—such as control charts, process monitoring techniques, and distance or divergence measures—as practical tools for detecting distribution shifts. It is essential to distinguish between statistically rigorous hypothesis tests, including both parametric and non-parametric approaches, and heuristic methods. Parametric tests, such as the t-test or F-test, rely on assumptions like normality and independence and typically require larger sample sizes; however, they offer precise Type I and Type II error control under these conditions. Non-parametric tests, such as the Kolmogorov–Smirnov test or the Friedman–Rafsky test, are equally valid in a statistical sense and provide robust inference without strong distributional assumptions. In contrast, heuristics like the Wasserstein Distance or Maximum Mean Discrepancy can be powerful in practice—especially in high-dimensional or complex settings—but often rely on empirically determined thresholds and lack formal guarantees on error rates. Table 1 summarizes these methods, and their detailed mathematical formulations are provided in Appendices D and E.
@@
There is no universally optimal method for detecting distributional shifts. Instead, Table 1 provides a toolbox for context-specific decision making. When assumptions are met, parametric tests offer high power and clean interpretability. For complex, high-dimensional, or weakly labeled post-deployment settings, non-parametric distributional distances—such as MMD, Wasserstein, and Friedman–Rafsky—are more robust and generalizable. Our appendix sections provide detailed mathematical formulations and guidelines for implementing these methods across a range of deployment environments in clinical AI systems.
Appendix C Turning Heuristics into Statistical Tests
{tblr}
cell11 = c=2, cell21 = r=3, cell22 = r=3, cell51 = r=4, cell52 = r=4, cell54 = c=2, cell64 = c=2, cell74 = c=2, cell84 = c=2, vlines, hline1-2,5,9 = -, hline3-4,6-8 = 3-5, hline5 = 2-, & Heuristics When to Use / Notes Data Distribution Assumptions
@@ Appendix C Turning Heuristics into Statistical Tests
M Shewhart Control Charts For process monitoring; detects sudden shifts normality, stable process, markov process
CUSUM Chart For detecting small, persistent shifts over time stable process, known target value
EWMA For detecting gradual changes with weighted historical data stable baseline, mean stationarity
Non-Param. Distr. Shifts Energy Distance Measures statistical distances between distributions
Wasserstein Distance When distributions have little or no overlap; captures shape/support shifts
Kullback-Leibler (KL) Divergence When distributions have complete support overlap; information-theoretic interpretation; asymmetric measure
Jensen-Shannon (JS) Divergence Bounded symmetric variant of KL divergence; use if distributions may not overlap; symmetric measure
@@
To convert the divergence measure into a test statistic with a known null distribution, we proceed as follows:
@@
        Split the permuted data into two groups of sizes $n_{0}$ and $n_{1}$.
@@
        Compute the divergence $T_{b}$ between these permuted groups.
@@
    Compute the p-value:
    $p=\frac{1+\sum_{b=1}^{B}\mathbf{1}\left(T_{b}\geq T_{\text{obs}}\right)}{1+B}$
@@
Appendix D Parametric Tests and Heuristics
D.1 Mean Shift
@@
where $\bar{X}_{k}$ and $s^{2}_{k}$ denote the sample mean and variance of group $k$, and $n_{k}$ represents the sample size, with degrees of freedom $df=n_{0}+n_{1}-2$. The null hypothesis ($H_{0}$) assumes no difference in means ($\mu_{0}=\mu_{1}$), while the alternative hypothesis ($H_{1}$) suggests a shift in mean. This test is appropriate for detecting mean shifts in $p_{t_{1}}$ when normality assumptions hold. With the critical value of
@@
$df=\frac{\left(\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}\right)^{2}}{\frac{\left(\frac{s_{1}^{2}}{n_{1}}\right)^{2}}{n_{1}-1}+\frac{\left(\frac{s_{2}^{2}}{n_{2}}\right)^{2}}{n_{2}-1}}$
@@
$\sigma^{2}_{E_{t}}$ $=\frac{\lambda}{2-\lambda}(1-(1-\lambda)^{2t})\sigma_{0}$
@@
where $\rho$ is a parameter. Note: we assume that varience remains unchanged after the distribution shift.
The univariate approaches discussed in above are computationally efficient, whoever they can miss complex feature interactions and face multiple testing challenges when monitoring many features simultaneously [94]. While all of the presented methods can be generalized to the multivariate setting, known as Multivariate Statistical Process (MSPC) control charts. The major limitation is the assumption that the processes distribution is multivariate normal [95], methods may not capture complex nonlinear relationships.
D.2 Variance Shift
@@
Appendix E Non-Parametric Tests and Heuristics
E.1 Mean Shift

### v3

@@ 6 Alternative Views
Several well–established research threads could serve as alternatives to the post-deployment monitoring problem. We examine the pros and cons of the top three alternatives to our proposed approach: continual learning, Bayesian change–point detection, and conformal‐prediction–based monitoring.
@@
or lifelong learning algorithms update model parameters online to accommodate non-stationary data distributions Ring (1994); Parisi et al. (2019); Chen and Liu (2018). Advantages include low latency adaptation as well as theoretical guarantees. Models can react to drift on the very next batch, which is attractive when label feedback is cheap. Additionally, PAC-Bayesian Alquier and Guedj (2016) or regret bounds are available for certain online update rules. However, this comes with several limitations: label requirement, auditability and traceability and silent failure. State-of-the-art continual learners still rely on frequent ground-truth labels to avoid catastrophic forgetting Rolnick et al. (2019). In clinical applications, those labels are costly or delayed. Regulatory guidelines (e.g. FDA SaMD) require reproducible model versions. Online updates create a moving target that complicates version control, performance analysis, and root-cause analysis U.S. Food and Drug Administration (2017). Finally, without an external test, online updates can chase noisy fluctuations and cause silent accuracy drops Mallick et al. (2022), which in turn may widen performance gaps for under-represented sub-groups in the data Yang et al. (2023).
@@
offers a probabilistic approach by maintaining a posterior distribution over run lengths and updating this belief as new data stream in Adams and MacKay (2007); Fearnhead and Liu (2007). This method provides coherent uncertainty quantification, allowing drift detection systems to trigger alarms based on posterior probabilities. It is also sequentially efficient: when using conjugate-exponential models, updates can be computed in constant amortized time per observation. However, BOCPD relies heavily on calibrated priors, which are rarely available or reliable in clinical contexts. Furthermore, applying BOCPD to high-dimensional data often requires approximate inference via particle filters or sequential Monte Carlo methods, which may be computationally infeasible for hospital-scale EHR feeds. Importantly, while BOCPD indicates when a change occurred, it does not identify where the change took place—limiting its usefulness in root-cause analysis and mitigation planning.
@@
offers a distribution-free framework for post-deployment monitoring. These methods maintain finite-sample validity under the assumption of exchangeability and have been extended to detect drift by monitoring conformal $p$-values or exchangeability martingales Vovk et al. (2022); Ho and Wechsler (2016); Johansson et al. (2022). They are attractive in that they do not require strong parametric assumptions, and variants such as Mondrian or conditional conformal predictors can operate effectively using weak labels. Still, this approach has notable weaknesses. The assumption of exchangeability is fragile in real-world settings like healthcare, where temporal, site-level, and treatment-based correlations are pervasive and violate i.i.d. conditions Barber et al. (2021). Furthermore, the signal provided by conformal methods is often blunt: they flag distributional shift only after prediction sets inflate, and do not distinguish between covariate and concept drift, nor do they identify the specific subgroups affected. As such, they tend to be reactive rather than proactive—triggering alarms after performance degrades, rather than before.
@@
Two-sample hypothesis testing combines rigorous error control, label efficiency, and interpretability. It provides explicit $\alpha$-level guarantees on type I error and power-based control of type II error. Unlike conformal methods, label-free tests such as kernel MMD Borgwardt et al. (2006) compare input distributions directly and highlight the feature regions driving the difference, providing regulator-friendly, interpretable evidence of shift. Detection is also modular: once drift is identified, retraining, recalibration, or continual learning can follow as appropriate. In contrast, continual learning, Bayesian monitoring, and conformal prediction each address part of the monitoring problem but fall short on one or more axes—supervision cost, statistical guarantees, or auditability. Hypothesis testing, by covering all three, is the most robust and regulator-ready foundation for monitoring AI systems in healthcare.
7 Limitations and Future Directions
This position paper focuses on statistically valid, label-efficient post-deployment monitoring and intentionally stops short of a fully causal treatment of failure attribution or label-semantics dynamics; we discuss monitoring of surrogate-label drift but leave causal analyses to future work. While we do not assume i.i.d. behavior across time (i.e., we assume that the data distribution has shifted relative to pre-deployment), we do assume that data are locally i.i.d. within pre- and post-deployment windows; extending the framework to handle explicit temporal dependence, time-series change detection, and broader nonstationarity is a natural next step.
While our formulation is modality-agnostic—operating on learned embeddings and applying the same drift, performance, and subgroup tests—we note a trade-off: higher dimensionality increases sample requirements and can reduce power. Consequently, label-efficient tactics (e.g., sequential looks and active labeling) together with clear reporting of margins and thresholds are key.
In the paper, we present statistical tests with binary outcomes. A natural extension is to treat validity as a continuum shaped by Pareto trade-offs among false-alarm control, detection delay, label budget, and subgroup granularity. Different operating points allocate limited risk and labeling resources differently across populations and time, tracing a frontier of feasible guarantees. This mirrors fairness–accuracy trade-offs in risk scoring: incompatible desiderata cannot be simultaneously optimized, so practitioners must select context-specific operating points with explicit, transparent priorities (Kleinberg et al., 2017).
While our work focuses on data- and model-level monitoring, deployments should also track downstream clinical impact—e.g., shifts in treatment patterns, workflow latency, and patient outcomes—using pragmatic designs with governance-backed thresholds; this operational layer is complementary to our scope. Finally, while we present a traditional ML formulation, emerging generative tools (e.g., LLMs, diffusion models, agents, and synthetic-data pipelines) introduce additional monitoring challenges—including prompt and data-provenance drift, stochastic output variability and hallucinations, content and usage safety, and generator–downstream feedback loops—that our framework does not yet cover. Extending statistically valid monitoring to these generative settings is an important direction for future work.
8 Conclusion
@@
Acknowledgments
This work was supported in part by the Office of Naval Research (ONR) under grant N00014-21-1-2615, the John and Tami Marick Foundation, and the National Science Foundation (NSF) under grant CCF-2048223.
Appendix A Reasons for the Data Shift
Reason Example
Covariate Shift Changes in Demographics Changes in patient demographics alter input features like age or comorbidities [66]
Data Quality Issues New EHR system leads to missing test results or erroneous data entry [66]
Changes in Data Collection Methods Transitioning to a new system changes lab result units or reference ranges [28]
Regulatory and Compliance Changes Privacy regulations limit the use of certain features critical for predictions [75]
Adversarial Attacks and Data Poisoning Falsifying patient records skews model predictions [24]
Ethical Constraints and Bias Mitigation Efforts Bias-mitigation interventions can change which features are used for prediction [68]
Changes in Measurement Techniques New lab assay for troponin provides results that aren’t comparable to the previous method [61]
Changes in Population Health Trends Aging populations or increases in chronic conditions shift the input feature distribution [66]
Label S Seasonal Disease Prevalence Flu prevalence increases in winter, changing the proportion of positive cases [38]
Seasonal Health Campaigns Public health campaigns can influence patient behavior [66]
Weather-related health issues Weather-related health issues can arise due to heatwaves or cold snaps [66]
Concept Drift Changes in Treatment Protocols New clinical guidelines alter the relationship between symptoms and outcomes [66]
Feedback Loops Additional care prevents readmission, and the model misinterprets this as reduced risk [1]
External Changes in Clinical Practice Adoption of new surgical techniques changes complication patterns [66]
Changes in Related Policies or Economic Factors Insurance policy changes alter patient behavior and readmission patterns [66]
Emergence of Unmeasured Confounding Variables Environmental hazards change disease presentation patterns [101]
Interaction Effects from Concurrent Models A new AI triage system changes patient flow and case severity [2]
Evolution of Disease Characteristics New variants alter symptom-outcome relationships [40]
Table 2: Reasons and examples for covariate shift, label shift and concept drift in the healthcare domain.
Appendix B Label Shift
Following the taxonomy in [64], label shift (also called prior probability shift or target shift) corresponds to the case where the marginal distribution of outcomes changes while the conditional distribution of features given outcomes remains fixed. Formally, label shift occurs when $p_{t_{1}}(y)\neq p_{t_{0}}(y)$ for some $y\in\mathcal{Y}$, while $p_{t_{1}}(\mathbf{s},\mathbf{c}\mid y)=p_{t_{0}}(\mathbf{s},\mathbf{c}\mid y)$ holds $\forall y\in\mathcal{Y}$ and $\forall(\mathbf{s},\mathbf{c})\in\mathcal{S}\times\mathcal{C}$. This type of shift is prevalent in the settings, where disease prevalence naturally fluctuates due to seasonal patterns, public health interventions, demographic changes, and evolving clinical practices.
To detect label shift, we formulate a hypothesis test on the marginal label distribution. Let $p_{t_{0}}(y)$ and $p_{t_{1}}(y)$ denote the marginal distributions of outcomes at $t_{0}$ and $t_{1}$. We test:
$H_{0}:$ $\quad p_{t_{0}}(y)=p_{t_{1}}(y)\quad\forall y\in\mathcal{Y}$ (10)
$H_{1}:$ $\quad p_{t_{0}}(y)\neq p_{t_{1}}(y)\quad\exists y\in\mathcal{Y}$
Note: we assume that $p_{t_{1}}(\mathbf{s},\mathbf{c}\mid y)=p_{t_{0}}(\mathbf{s},\mathbf{c}\mid y)$ holds $\forall y\in\mathcal{Y}$ and $\forall(\mathbf{s},\mathbf{c})\in\mathcal{S}\times\mathcal{C}$. When label shift is detected, several mitigation strategies are available that do not require full model retraining [84]. For probabilistic classifiers, output probabilities can be adjusted using the ratio $\frac{p_{t_{1}}(y)}{p_{t_{0}}(y)}$ to account for changed base rates, while binary classifiers can undergo decision threshold recalibration to maintain desired operating characteristics (e.g., sensitivity/specificity trade-offs) under the new class distribution. When retraining is feasible, importance weights $w_{i}=\frac{p_{t_{1}}(y_{i})}{p_{t_{0}}(y_{i})}$ can be applied to training samples to simulate the target distribution [54]. Additionally, Expectation-Maximization approaches such as Black Box Shift Estimation (BBSE) [54] can estimate label shift and correct predictions using only unlabeled post-deployment data and a confusion matrix, though these methods carry their own assumptions and limitations.
OP: Label Scarcity
Detecting label shift requires access to ground truth labels $Y$ at both pre- and post-deployment stages to estimate $p_{t_{0}}(y)$ and $p_{t_{1}}(y)$. While label shift detection requires fewer labels than concept drift detection (only marginal distributions rather than joint distributions), the fundamental label scarcity problem remains. Active sampling strategies that prioritize diverse samples across the feature space (rather than focusing on model uncertainty as in active learning) may be more appropriate for label shift detection. Stratified sampling approaches that ensure representation across key demographic and clinical subgroups can improve prevalence estimation efficiency.
Appendix C Performance Metrics
@@
Appendix D Choosing A Statistical Hypothesis Test and Heuristics
@@
In addition to formal test statistics, our framework includes heuristics—such as control charts, process monitoring techniques, and distance or divergence measures—as practical tools for detecting distribution shifts. It is essential to distinguish between statistically rigorous hypothesis tests, including both parametric and non-parametric approaches, and heuristic methods. Parametric tests, such as the t-test or F-test, rely on assumptions like normality and independence and typically require larger sample sizes; however, they offer precise Type I and Type II error control under these conditions. Non-parametric tests, such as the Kolmogorov–Smirnov test or the Friedman–Rafsky test, are equally valid in a statistical sense and provide robust inference without strong distributional assumptions. In contrast, heuristics like the Wasserstein Distance or Maximum Mean Discrepancy can be powerful in practice—especially in high-dimensional or complex settings—but often rely on empirically determined thresholds and lack formal guarantees on error rates. Table 1 summarizes these methods, and their detailed mathematical formulations are provided in Appendices F and G.
@@
There is no universally optimal method for detecting distributional shifts. Instead, Table 1 provides a toolbox for context-specific decision making. When assumptions are met, parametric tests offer high power and clean interpretability. For complex, high-dimensional, or weakly labeled post-deployment settings, non-parametric distributional distances—such as MMD, Wasserstein, and Friedman–Rafsky—are more robust and generalizable. Refer to Appendices F and G for the descriptions of the parametric and non-parametric tests, respectively.
Appendix E Turning Heuristics into Statistical Tests
Heuristics When to Use / Notes Data Distribution Assumptions
@@ Appendix E Turning Heuristics into Statistical Tests
M
F.1 Shewhart Control Charts For process monitoring; detects sudden shifts normality, stable process, markov process
F.1 CUSUM Chart For detecting small, persistent shifts over time stable process, known target value
F.1 EWMA For detecting gradual changes with weighted historical data stable baseline, mean stationarity
Non-Param.
Distr. Shifts
G.3 Energy Distance Measures statistical distances between distributions
G.3 Wasserstein Distance When distributions have little or no overlap; captures shape/support shifts
G.3 Kullback-Leibler (KL) Divergence When distributions have complete support overlap; information-theoretic interpretation; asymmetric measure
G.3 Jensen-Shannon (JS) Divergence Bounded symmetric variant of KL divergence; use if distributions may not overlap; symmetric measure
@@ G.3 Wasserstein Distance When distributions have little or no overlap; captures shape/support shifts
The permutation test simulates the distribution of a test statistic (e.g., Jensen–Shannon divergence) under the null hypothesis $H_{0}$, and computes the $p$-value using this null distribution based on the observed two samples, $\mathcal{D}_{t_{0}}$ and $\mathcal{D}_{t_{1}}$, to determine whether to reject $H_{0}$. The permutation test is performed as follows.
@@
        Split the permuted data into two groups of sizes $n_{0}$ and $n_{1}$ according to the permuted labels.
@@
        Compute the test statistic $T_{b}$ using (1) based on the two permuted groups.
@@
    Compute the $p$-value (right-sided) [^5] Left-sided or two-sided $p$-values can be computed analogously without loss of generality.:
    $p=\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}\left(T_{b}\geq T_{\text{obs}}\right).$
@@
Appendix F Parametric Tests and Heuristics
F.1 Mean Shift
@@
where $\bar{X}_{k}$ and $s^{2}_{k}$ denote the sample mean and variance of group $k$, and $n_{k}$ represents the sample size, with degrees of freedom $df=n_{0}+n_{1}-2$. The null hypothesis ($H_{0}$) assumes no difference in means ($\mu_{0}=\mu_{1}$), while the alternative hypothesis ($H_{1}$) suggests a shift in mean. This test is appropriate for detecting mean shifts in $p_{t_{1}}$ (the post-deployment distribution) when normality assumptions hold. With the critical value of
@@
$df=\frac{\left(\frac{s_{0}^{2}}{n_{0}}+\frac{s_{1}^{2}}{n_{1}}\right)^{2}}{\frac{\left(\frac{s_{0}^{2}}{n_{0}}\right)^{2}}{n_{0}-1}+\frac{\left(\frac{s_{1}^{2}}{n_{1}}\right)^{2}}{n_{1}-1}}$
@@
$\sigma^{2}_{E_{t}}$ $=\frac{\lambda}{2-\lambda}(1-(1-\lambda)^{2t})\sigma^{2}_{0}$
@@
where $\rho$ is a parameter. Note: we assume that variance remains unchanged after the distribution shift.
The univariate approaches discussed above are computationally efficient, but they can miss complex feature interactions and face multiple testing challenges when monitoring many features simultaneously [63]. While all of the presented methods can be generalized to the multivariate setting, known as Multivariate Statistical Process (MSPC) control charts. The major limitation is the assumption that the processes distribution is multivariate normal [76], methods may not capture complex nonlinear relationships.
F.2 Variance Shift
@@
Appendix G Non-Parametric Tests and Heuristics
G.1 Mean Shift